OpenStudy (anonymous):

Find the exact value in radians. Super bad at radians so an explanation would be nice as well since I have quite a few of these problems!The test is tomorrow!!! Also no calculator is allowed.

OpenStudy (anonymous):

I'm also on my phone so I apologize for any delay of reply.

OpenStudy (anonymous):

A radian is a unit of measure for angles which usually corresponds to the arc length of the angle measured

OpenStudy (anonymous):

\[1 radian= \frac{ 180 degrees }{ \pi } \approx 57.2958 degrees\]

OpenStudy (anonymous):

I've dealt with radians in class it's just that I'm terrible at them and I dont remember how to solve for them with these particular type of problems

OpenStudy (anonymous):

This conversion factor can help you solve for radians and degrees and back\[degrees = radians \times \frac{ 180 degrees }{ \pi }\] \[radians = degrees \times \frac{ \pi }{ 180 degrees }\]

OpenStudy (anonymous):

Okay. I've tried that before, maybe I'm just doing my mental math wrong. So is it ok for me to solve in degrees then just convert?